# Interact-RAG: Reason and Interact with the Corpus, Beyond Black-Box Retrieval

### v1

@@ 5.1 Retrieval-augmented Generation
Retrieval-Augmented Generation (RAG) is a prevailing method to enhance LLMs with external information (Lewis et al., 2020). Basic RAG relies on static embedding-based retrieval, which may suffer from information omission (Gao et al., 2023). To address this, various studies propose tree-based or graph-based index to improve retrieval robustness (Jin et al., 2025b; Edge et al., 2024; Luo et al., 2025). Another direction focuses on improving the retrieval pipeline. Iterative RAGs were introduced to progressively refine information through multi-step retrieval (Trivedi et al., 2023; Chan et al., 2024; Hui et al., 2025). Recent agentic methods provide more flexibility, where the LLM autonomously orchestrate the entire RAG pipeline (Gao et al., 2025). Methods such as MA-RAG (Nguyen et al., 2025), Search-O1 (Li et al., 2025b), and MCTS-RAG (Hu et al., 2025) implement prompt-driven strategies, leveraging multiple agentic modules. End-to-end approaches like Search-R1 (Jin et al., 2025a), InForage (Qian & Liu, 2025), and SimpleDeepSearcher (Sun et al., 2025) adopt SFT and RL to create fully autonomous agents. Despite the effectiveness of above approaches, they often operate within a black-box retrieval paradigm, limiting the analysis and control. Addressing this, our work explores an interactive framework with fine-grained retrieval manipulation, supporting improved reasoning and adaptability.
@@ 5.2 Reasoning-Enhanced LLM Agent
Enhancing LLMs with reasoning has become a prevailing research focus (Xu et al., 2025). The strategies span prompting-based approaches like Chain-of-Thought Wei et al. (2022), and training-optimized models like OpenAI o1/o3/o4 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025). To support broader scenarios, various works leverage reasoning to improve the performance of LLM agents(Ferrag et al., 2025), training them to use tools and solve complex problems (Lu et al., 2025b; Shen et al., 2025). They explore various dimensions, including the construction of high-quality training data (Li et al., 2025a; Shi et al., 2025), the refinement of reward signals (Zhao et al., 2025; Qian & Liu, 2025), and the optimization of reinforcement learning algorithms (Dong et al., 2025; Lu et al., 2025a). While these works have made great advances concentrating on the agent’s training, our focus is distinct: we redesign the interaction paradigm for RAG agents and leverage the reasoning capability to enable fine-grained manipulation.
@@ Appendix A The Use of LLMs
Appendix B Additional Experiments and Demonstration
B.1 Additional Experimental Results
Method Multi-Hop QA Single-Hop QA AVG
HotpotQA 2Wiki. Musique Bamboogle NQ PopQA
EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1
Interact-RAG-2.5 47.8 61.6 63.6 71.0 30.9 39.5 47.6 61.1 43.7 52.9 51.6 54.4 47.5 56.8
Table 4: Additional results in Exact Match (EM) and F1 scores across various benchmarks. Bold and underline denote the best and second-best performance respectively. All models are based on the instruction-tuned Qwen2.5-7B backbone. Our Interact-RAG (Qwen2.5 version) achieves a relative improvement of 10.1% over the advanced baseline Search-R1. Notably, our Interact-RAG was trained on 12K QA data, a small fraction of the 170K QA pairs used for Search-R1. This data disparity also explains Search-R1’s stronger performance on the NQ dataset.
B.2 Case Study
@@
Appendix C Additional Implementation Details
C.1 Dataset Details
@@
C.2 Implementation of our Corpus Interaction Engine
@@
The core functionalities are realized as follows: (1) For Semantic-Search, we implemented a retriever based on the e5-base-v2 model (Wang et al., 2022), using the prevailing ChromaDB (core team, 2025) as our underlying vector database. (2) Exact-Search is built upon the Full-Text Search (FTS) module of SQLite database (Bhosale et al., 2015), which returns results ranked by the BM25 scores between query keywords and text chunks. (3) In the Fusion Stage of semantic and exact search, we first normalize the scores of the top-20 chunks from each search strategy. These scores are then aggregated via a weighted sum, according to the weight specified by the agent, and the high-scoring chunks are ultimately returned. (4) For Entity-Matching, also based on SQLite FTS, the engine precisely locates text segments containing a specified term and appends the three most query-related snippets to the result. (5) Simpler actions like Include-Docs and Exclude-Docs are handled directly through basic filtering operations.
C.3 Experimental Details
To ensure generality and alignment, our action pipeline is implemented using the official reasoning and tool-use template from Qwen3 (Yang et al., 2025), which inherently utilizes <think>, <tool_call>, and <tool_response>  tags. In the Supervised Fine-Tuning (SFT) stage, we employ the Llama-Factory framework (Zheng et al., 2024), training for 2 epochs with a learning rate of $2\times 10^{-5}$ and a batch size of 128. Following this, the agent is refined through Reinforcement Learning (RL) using the verl framework (Sheng et al., 2025). The RL phase involves multi-turn agent training for 2 epochs, with a policy learning rate of $1\times 10^{-6}$, a batch size of 128, a maximum of 7 interaction turns, and the rollout-num of 8. All experiments were conducted on a cluster of 8 NVIDIA A100 (80GB) GPUs.
@@ C.4 LLM Prompts
End-to-End Prompt for Interactive RAG Agent You are a strategic AI research assistant. Your task is to answer user questions by leveraging a search tool. You must operate in a systematic, iterative loop of planning, acting, and analyzing. ## Your Research Process ### 1. Understand & Plan: Understand the user’s question and create a search plan. - First, thoroughly analyze the user’s question. Identify key concepts, entities, and any constraints. - If the question is straightforward, formulate a single, comprehensive search query that is most likely to yield the final answer. - If the question is complex, break it down and define the clear and specific sub-tasks. Please outline the sub-questions and desired outcomes for each step. - If some sub-tasks can be executed parallelly, you should point out it. - You should first perform thinking, and output the primary plan in the list format. ### 2. Execute the Search Based on the current state and previous analysis, call the execute_search_plan tool to perform the search. - The parameter semantic_query is primary and required. It should be a clear and concise query to search the needed information. - There are also serval optional parameters to improve the search results. You should perform analysis and adapt the parameters actively and reasonably. ### 3. Observe & Iterate Analyze the retrieved context, and decide the next action. - If the received context is not good, you should reflect to improve the search, and execute the search tool again with the refined parameters. - If you get sufficient information for the sub-question, you can proceed to the next sub-task with another search execution. You should make sure the search for current step is enough, don’t be overly confident about some noise. - If you have gathered sufficient evidence to construct a complete answer for the whole question, you should conclude the final answer with no more function-calls. - You don’t need to follow the primary search plan strictly. You can adapt your strategy based on the retrieved context and your analysis. ## Final Answer Formulation Once you have enough evidence to get the final answer, you can just conclude it. The final answer must be concise and direct words.
@@
Prompt for the Executor within our Workflow You are a specialized searching execution agent. You will be presented with a user’s query and prior search results with analysis. Your sole purpose is to perform one of two specific actions: either call the execute_search_plan tool or provide the final answer. # # Your Actions (Choose ONE): 1. Execute a Search: - Based on the provided instructive analysis, you should identify the proper query and parameters. - Your task is to call the tool with the appropriate parameters. Note: The semantic_query parameter is required. It should be clear and specific. If the previous instructions do not provide a query, you should formulate one. There are also serval optional parameters to refine search results. You can make up to 2 seperate calls in one turn, if needed (i.e., some sub-tasks can be executed parallelly). 2. Formulate the Final Answer: Based on the provided information and the analysis, if the evidence is sufficient to answer the user’s whole original question, you should provide a final answer. The final answer must be concise and direct words

### v3

@@ 5.1 Retrieval-augmented Generation
Retrieval-Augmented Generation (RAG) is a prevailing method to enhance LLMs with external information (Lewis et al., 2020). Basic RAG relies on static embedding-based retrieval (Lewis et al., 2020; Li et al., 2025e; d), which may suffer from information omission (Gao et al., 2023). To address this, various studies propose tree-based or graph-based index to improve retrieval robustness (Jin et al., 2025b; Edge et al., 2024; Luo et al., 2025). Another direction focuses on improving the retrieval pipeline. Iterative RAGs were introduced to progressively refine information through multi-step retrieval (Trivedi et al., 2023; Chan et al., 2024; Hui et al., 2025). Recent agentic methods provide more flexibility, where the LLM autonomously orchestrates the entire RAG pipeline (Gao et al., 2025; Wan et al., 2025). Methods such as MA-RAG (Nguyen et al., 2025), Search-O1 (Li et al., 2025b), and MCTS-RAG (Hu et al., 2025) implement prompt-driven strategies, leveraging multiple agentic modules. End-to-end approaches like Search-R1 (Jin et al., 2025a), InForage (Qian and Liu, 2025), and SimpleDeepSearcher (Sun et al., 2025) adopt SFT and RL to create fully autonomous agents. Despite the effectiveness of above approaches, they often operate within a black-box retrieval paradigm, limiting the analysis and control. Addressing this, our work explores an interactive framework with fine-grained retrieval manipulation, supporting improved reasoning and adaptability.
@@ 5.2 Reasoning-Enhanced LLM Agent
Enhancing LLMs with reasoning has become a prevailing research focus (Xu et al., 2025). The strategies span prompting-based approaches like Chain-of-Thought Wei et al. (2022), and training-optimized models like OpenAI o1/o3/o4 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025). To support broader scenarios, various works leverage reasoning to improve the performance of LLM agents(Ferrag et al., 2025), training them to use tools and solve complex problems (Lu et al., 2025b; Shen et al., 2025). They explore various dimensions, including the construction of high-quality training data (Li et al., 2025a; Shi et al., 2025), the refinement of reward signals (Zhao et al., 2025; Qian and Liu, 2025), and the optimization of reinforcement learning algorithms (Dong et al., 2025; Lu et al., 2025a). While these works have made great advances concentrating on the agent’s training, our focus is distinct: we redesign the interaction paradigm for RAG agents and leverage the reasoning capability to enable fine-grained manipulation.
@@ 6 Conclusion
Ethics statement
This work adheres to the ethical guidelines set forth by ICLR 2026. We have conducted our research with a commitment to avoiding harm, ensuring honesty and transparency in our methodology and reporting. We have made concerted efforts to identify and mitigate potential biases in our data and algorithms to ensure fairness. Furthermore, our research respects individual privacy, and we have complied with all applicable regulations and ethical standards regarding data use.
Reproducibility statement
In this work, we present Interact-RAG, a new paradigm empowering LLM agents with fine-grained control over the information-seeking process. To ensure reproducibility and facilitate further research, we provide the source code in the supplementary materials. Additionally, we offer comprehensive documentation in Section 4.1 and Appendix D, covering dataset details, training parameters, environment constructions, and experimental configurations.
@@ Appendix A The Use of LLMs
Appendix B Limitations
Despite the effectiveness of Interact-RAG, we acknowledge two limitations: (1) RL Optimization: Our RL stage currently relies on outcome-based rewards. While this aligns with some established GRPO practices, incorporating granular process rewards or advanced RL algorithm could further enhance learning efficiency. (2) Test-time Cost: Our explicit Chain-of-Thought reasoning trajectory incurs higher latency and token usage, compared to non-reasoning modes. This test-time scaling reflects the trade-off inherent in reasoning models, where increased compute is exchanged for superior accuracy. In future work, we will explore more advanced RL designs and optimize the cost-effectiveness.
Appendix C Additional Demonstration and Experiments
C.1 Case Study
@@
C.2 Generalization to other Domains
For more comprehensive evaluation over other non-wiki domains, we conducted additional experiments on the MultiHop-RAG benchmark (Tang and Yang, 2024). Its corpus consists of news covering diverse fields such as technology, health, and business. As shown in the Table 5, Interact-RAG significantly outperforms baselines in this out-of-distribution setting. This further demonstrates the effectiveness and robustness of our interaction-reasoning paradigm.
Method EM F1
7B Models
Std-RAG-7B 57.5 59.4
MA-RAG-7B 50.2 52.3
Search-R1-7B 73.0 74.1
Interact-RAG-7B 79.4 80.4
8B Models
Std-RAG-8B 63.4 65.1
MA-RAG-8B 66.3 68.7
Interact-RAG-8B 81.4 82.3
Table 5: Performance results on the MultiHop-RAG benchmark (non-wiki domain).
C.3 Retrieval Quality
To assess retrieval quality beyond final answer accuracy, we evaluate the coverage of supporting facts across three multi-hop benchmarks, since they provide human-annotated evidence. We calculated the proportion of supporting facts covered within the retrieved context. As shown in Table 6, our Interact-RAG consistently outperforms other baselines. This confirms that our active interaction paradigm effectively improves the information retrieval and gathering process.
Method 2Wiki Musique HotpotQA
Standard-Retrieval 0.434 0.127 0.502
MA-RAG-7B 0.620 0.256 0.573
MA-RAG-8B 0.631 0.277 0.581
Search-R1-7B 0.608 0.256 0.609
Interact-RAG-7B 0.686 0.316 0.632
Interact-RAG-8B 0.706 0.335 0.637
Table 6: Retrieval quality evaluation. We report the coverage of supporting facts within the retrieved context.
C.4 Cost-Performance Trade-off Analysis.
Table 7 provides a detailed comparison of computational costs on the 2WikiMultiHopQA dataset. The evaluation was conducted on the NVIDIA A100 GPU utilizing the vLLM inference engine, and the results are averaged per question. To ensure the consistent comparison, the methods are all based on the Qwen2.5-7B model. We acknowledge that Interact-RAG incurs higher latency and token consumption compared to standard baselines. This increase stems from the explicit reasoning trajectory. However, this design aligns with the emerging paradigm of reasoning models and test-time scaling (e.g., OpenAI-o1 (Jaech et al., 2024), DeepSeek-R1 (Guo et al., 2025)), where increased inference compute is a necessary trade-off for superior accuracy over complex tasks.
Furthermore, we conducted a control experiment applying a Best-of-N ($N=3$) sampling strategy to the strong baseline, Search-R1. While this approach increased the computational cost by over $3\times$, the performance only improved from 50.9 to 56.2, still falling short of ours. This demonstrates the effectiveness of Interact-RAG.
Method EM Score Latency (s) Tool Time (s) Out Tok. In Tok.
MA-RAG-7B 39.6 15.1 5.09 181 3812
Search-R1-7B 50.9 17.2 5.61 192 4289
Interact-RAG-7B 63.6 36.4 4.67 713 5216
Table 7: Detailed computational cost analysis on the 2WikiMultiHopQA dataset. We report the Exact Match (EM) score alongside average wall-clock latency, tool execution time, and accumulated token consumption per query.
C.5 Fusion Policy for Multi-faceted Retrieval
In the multi-faceted retrieval, we need to fuse the results from two distinct strategies. In our implementation, we apply min-max normalization to both scores, and then perform the weighted aggregation, where the weight is assigned by the LLM agent. Now, we compare our fusion policy with two standard rank fusion methods, including Reciprocal Rank Fusion (RRF) and CombMNZ (Cormack et al., 2009). As shown in Table 8 , our implementation yields performance comparable or slightly superior. And we would like to clarify that the fusion mechanism is a small functional design in our interaction engine, so its variations do not significantly alter the overall performance.
Fusion Method 2Wiki Musique PopQA
RRF 68.7 34.6 55.3
CombMNZ 69.8 32.9 56.0
Interact-RAG 69.6 34.8 56.0
Table 8: Comparison of different fusion strategies (reported in EM Scores).
C.6 Overhead of our Corpus Interaction Engine
For the memory footprint, on a corpus of 21.8K documents (280K chunks), the standard vector database (based on Chromadb (team, 2025)) occupies 2.3GB. And our additional interaction component (based on SQLite FTS-based (SQLite, 2025)) occupies only 0.3GB, representing a negligible memory addition. Besides, we evaluate the deployment performance of our engine using the 2Wiki dataset, in an 8-core CPU environment with 96 concurrent request clients. For the wall-clock time, our interaction processing latency is 1.82s per iteration, only $\sim 3\%$ higher than the standard vector-based retrieval (1.76s), which is negligible.
Appendix D Additional Implementation Details
D.1 Dataset Details
@@
D.2 More details of Corpus Interaction
@@
The core functionalities are realized as follows: (1) For Semantic-Search, we implemented a retriever based on the e5-base-v2 model (Wang et al., 2022), using the prevailing ChromaDB (team, 2025) as our underlying vector database. (2) Exact-Search is built upon the Full-Text Search (FTS) module of SQLite database (Bhosale et al., 2015), which returns results ranked by the BM25 scores between query keywords and text chunks. (3) In the Fusion Stage of semantic and exact search, we first apply the min-max normalization over the scores of the top-20 chunks from each search strategy. These scores are then aggregated via a weighted sum, according to the weight specified by the agent, and the high-scoring chunks are ultimately returned. (4) For Entity-Matching, the LLM agent is responsible and capable to provide the specific entity keywords, based on the query and history context. Then, we utilize SQLite FTS to retrieve sentences containing the entity terms, where the words are normalized for comparison. After that, we retain and append the three most relevant small sentences based on the current sub-query. (5) Simpler actions like Include-Docs and Exclude-Docs are handled directly through basic filtering operations. Regarding filter persistence, the LLM is instructed to manage the state by reissuing filter instructions if needed.
The engine follows a deterministic pipeline to resolve concurrent actions: (1) Retrieval scores are fused to form a candidate list. (2) Doc-filters (include/exclude) are applied to explicitly retain or remove chunks. (3) Top-ranked chunks are selected according to the chunk budget. (4) Unique entity-specific short snippets are appended.
It is worth noting that while the agent may invoke multiple strategies in a single iteration, our engine avoids generating multiple large contexts. Instead, it produces a single consolidated context by aggregating these primitives. Specifically, multi-strategy retrieval yields a unified chunk set via score fusion, and Entity Match appends only concise snippets with negligible addition. While context-shaping may dynamically change the number of chunks, our ablation studies (Table 3) demonstrate that Interact-RAG still maintains strong performance without this module. Collectively, these suggest that our performance gains stem from fine-grained interaction strategies, rather than simply inflating the retrieval volume.
D.3 Experimental Details
All our experiments were conducted on a cluster of 8 NVIDIA A100 (80GB) GPUs. To ensure generality and alignment, our action pipeline is implemented using the official reasoning and tool-use template from Qwen3 (Yang et al., 2025), which inherently utilizes <think>, <tool_call>, and <tool_response>  tags. In the Supervised Fine-Tuning (SFT) stage, we employ the Llama-Factory framework (Zheng et al., 2024), training for 2 epochs with a learning rate of $2\times 10^{-5}$ and a batch size of 128. Following this, the agent is refined through Reinforcement Learning (RL) using the verl framework (Sheng et al., 2025). The RL phase involves multi-turn agent training for 2 epochs, with a policy learning rate of $1\times 10^{-6}$, a batch size of 128, a maximum of 7 interaction turns, and the rollout-num of 8. During the RL training, we observed the format error rate of 2% at the beginning. And this rate declined to 1.1% at step 50 and dropped to 0.04% by the final step (i.e., step-110), which also demonstrates the effectiveness of our RL optimization.
@@ D.4 LLM Prompts
End-to-End Prompt for Interactive RAG Agent You are a strategic AI research assistant. Your task is to answer user questions by leveraging a search tool. You must operate in a systematic, iterative loop of planning, acting, and analyzing. ## Your Research Process ### 1. Understand & Plan: Understand the user’s question and create a search plan. - First, thoroughly analyze the user’s question. Identify key concepts, entities, and any constraints. - If the question is straightforward, formulate a single, comprehensive search query that is most likely to yield the final answer. - If the question is complex, break it down and define the clear and specific sub-tasks. Please outline the sub-questions and desired outcomes for each step. - If some sub-tasks can be executed parallelly, you should point out it. - You should first perform thinking, and output the primary plan in the list format. ### 2. Execute the Search Based on the current state and previous analysis, call the execute_search_plan tool to perform the search. - The parameter semantic_query is primary and required. It should be a clear and concise query to search the needed information. - There are also several optional parameters to improve the search results. You should perform analysis and adapt the parameters actively and reasonably. ### 3. Observe & Iterate Analyze the retrieved context, and decide the next action. - If the received context is not good, you should reflect to improve the search, and execute the search tool again with the refined parameters. - If you get sufficient information for the sub-question, you can proceed to the next sub-task with another search execution. You should make sure the search for current step is enough, don’t be overly confident about some noise. - If you have gathered sufficient evidence to construct a complete answer for the whole question, you should conclude the final answer with no more function-calls. - You don’t need to follow the primary search plan strictly. You can adapt your strategy based on the retrieved context and your analysis. ## Final Answer Formulation Once you have enough evidence to get the final answer, you can just conclude it. The final answer must be concise and direct words.
@@
Prompt for the Executor within our Workflow You are a specialized searching execution agent. You will be presented with a user’s query and prior search results with analysis. Your sole purpose is to perform one of two specific actions: either call the execute_search_plan tool or provide the final answer. # # Your Actions (Choose ONE): 1. Execute a Search: - Based on the provided instructive analysis, you should identify the proper query and parameters. - Your task is to call the tool with the appropriate parameters. Note: The semantic_query parameter is required. It should be clear and specific. If the previous instructions do not provide a query, you should formulate one. There are also several optional parameters to refine search results. You can make up to 2 seperate calls in one turn, if needed (i.e., some sub-tasks can be executed parallelly). 2. Formulate the Final Answer: Based on the provided information and the analysis, if the evidence is sufficient to answer the user’s whole original question, you should provide a final answer. The final answer must be concise and direct words